OpenStudy (anonymous):

What does it mean by find all real values of x such that f(x) = 0? I don't really know how to apply that to this problem f(x)=12- (x^2) / 5

OpenStudy (anonymous):

it means set it equal to zero and solve

OpenStudy (anonymous):

graph it. Then look at where x = 0. then use your finger and drag it to the graph to see the Y values

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that's what I would do :D

OpenStudy (anonymous):

is it \[f(x)=12-\frac{x^2}{5}\]?

OpenStudy (anonymous):

noo 12 is with -x^2 sorry if I made it unreadable

OpenStudy (anonymous):

i got x^2=7

OpenStudy (anonymous):

or would it be sqrt +- 7?

OpenStudy (anonymous):

so \[\frac{12-x^2}{5}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then \[x=\pm\sqrt{12}=\pm 2\sqrt{3}\]

OpenStudy (anonymous):

Wait, what happened to the 5?

OpenStudy (anonymous):

you multiply through by 5

OpenStudy (anonymous):

so you get 60-5x^2 = 0

OpenStudy (anonymous):

a fraction is zero means the numerator is 0 you can ignore the 5

OpenStudy (anonymous):

\[\frac{x-2}{72}=0\] same as \[x-2=0\]

OpenStudy (anonymous):

ok wait, it says f(x) equals zero So why wouldn't it be just 12/5

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it say \[f(x)=0\] right, and you function is \[f(x)=\frac{12-x^2}{5}\]

OpenStudy (anonymous):

find all real values

OpenStudy (anonymous):

it is not asking for \[f(0)\] it is saying that \[f(x)=0\]

OpenStudy (anonymous):

you confused the two. \[f(0)=\frac{12-0^2}{5}=\frac{12}{5}\]

OpenStudy (anonymous):

but it is not saying that the input is zero. it is saying that the output is 0

OpenStudy (anonymous):

\[f(x)=0=\frac{12-x^2}{5}\]

OpenStudy (anonymous):

ok , so since its equal to zero you can ignore the denominator?

OpenStudy (anonymous):

yes because it is a number

OpenStudy (anonymous):

So, if it said to solve for x you would do the same correct?